# Spectral representations of interpolation spaces of reproducing kernel Hilbert spaces

### v1

@@ Abstract
In statistical learning theory, interpolation spaces of the form $\mathrm{L}^{2},H]_{\theta,r}$, where $H$ is a reproducing kernel Hilbert space, are in widespread use. So far, however, they are only well understood for fine index $r=2$. We generalise existing results from $r=2$ to all possible values of $r$. In particular, we present a spectral decomposition of such spaces, analyse their embedding properties, and describe connections to the theory of Banach spaces of functions. We additionally present exemplary applications of our results to regularisation error estimation in statistical learning.
@@ 1 Introduction
Since their introduction, reproducing kernel Hilbert spaces (RKHS) have turned out as fundamental objects in various branches of mathematical theory as well as in applications. As they combine the geometry of a Hilbert space with the continuous evaluation functionals of a function space, they posses many desirable properties. Apart from applications in functional analysis and numerics, statistical learning theory takes advantage of this situation by considering a large variety of kernel-based algorithms.
@@
are important for the analysis of kernel-based least squares regression, see e.g. [5], for estimations of the conditional mean embedding, see e.g. [11], and the analysis of learning curves, see e.g. [10]. Similarly, these spaces occur naturally for certain questions related to Gaussian processes, see e.g. [15], [9].
The main reason for choosing $2$ as the fine index in (1) is the spectral representation of these interpolation spaces. More precisely, the spaces (1) are given as the range of powers of the integral operator associated to the RKHS [18, Section 4]. Using eigenvalues and eigenfunctions of this operator hence yields many insights into interpolation spaces and turns out as very helpful in statistical learning theory.
@@
  • Fine index $1$ is of interest as it is allows for certain embedding that are not possible for other fine indexes, see e.g. [17, p. 4]
  • Fine index $\infty$ is of interest as it allows for a precise analysis of approximation errors, see e.g. [14, Theorem 3.1].
@@
that is, we allow for all possible fine indexes $1\leq r\leq\infty$. Our aim is to present methods and results that treat such interpolation spaces for all fine indexes $r$ in a joint fashion. The results, we will derive address the following main questions:
@@
  • The interpolation spaces consist of equivalence classes of functions at first. When can we identify the spaces (2) with Banach spaces consisting of functions?
  • When are the spaces (2) continuously embedded into $\mathrm{L}^{\infty}$?
The first question is clearly motivated by the practicality of spectral representations, as indicated by the papers mentioned earlier. Positive answers to the second question have applications in all situations where a point evaluation of functions is necessary. A main example is given by the theory of Gaussian processes in the context of machine learning, as paths cannot be equivalence classes but need to be defined everywhere, see e.g. [12, Assumption 1]. The third question will turn out to be tightly connected to the second one, but applications go beyond that. In particular, the statistical analysis of kernel-based learning algorithms relies on such embeddings, see e.g. [16, Theorem 7.23] [5, Section 3] [17, p. 4].
This work is organised as follows: We start by introducing the necessary notation and fundamental definitions in Section 2. Then the main results are provided in Section 3. We present an exemplary application to error estimation in statistical learning theory in Section 4. All proofs and the ideas behind the results will be given in Section 5. Finally, some auxiliary results can be found in two appendices.
@@ 2 Preliminaries
Given normed $\mathbb{R}$-vector spaces $E_{1}$ and $E_{2}$, we use $E_{1}\simeq E_{2}$ to denote isometric isomorphy, $E_{1}\triangleq E_{2}$ for norm equivalent spaces and $E_{1}\hookrightarrow E_{2}$ if $E_{1}\subset E_{2}$ and the embedding is continuous.
For a measurable space $(X,\mathcal{A})$, we denote the set of measurable functions $X\rightarrow\mathbb{R}$ by $\mathcal{L}^{0}(X)$. The set of bounded functions in $\mathcal{L}^{0}(X)$ is denoted by $\mathcal{L}^{\infty}(X)$ and this space is equipped with the supremum norm.
@@
Let us introduce some background from the theory of interpolation spaces of the real interpolation method. For more details, see [3, Chapter 5] or [4, Chapter 3]. In the following, we only consider interpolation spaces of Banach spaces $E_{1}$ and $E_{2}$ where $E_{1}\hookrightarrow E_{2}$.
To define the real interpolation spaces, we use the so-called $K$-method. To this end, for all $x\in E_{2}$ and $t>0$, we define the $K$-functional as
$K(x,t,E_{2},E_{1}):=\inf_{y\in E_{1}}\left(\|x-y\|_{E_{2}}+t\|y\|_{E_{1}}\right)\text{.}$ (3)
For $x\in E_{2}$ and the parameters $0<\theta<1$ and $1\leq r<\infty$, this enables us to set
$\|x\|_{[E_{2},E_{1}]_{\theta,r}}:=\left(\int_{0}^{\infty}t^{-\theta r-1}K^{r}(x,t,E_{2},E_{1})\mathrm{d}t\right)^{\frac{1}{r}}$ (4)
@@
$\|x\|_{[E_{2},E_{1}]_{\theta,\infty}}:=\sup_{t>0}\left(t^{-\theta}K(x,t,E_{2},E_{1})\right)\text{.}$ (5)
@@
$[E_{2},E_{1}]_{\theta,r}:=\{x\in E_{2}\,:\,\|x\|_{[E_{2},E_{1}]_{\theta,r}}<\infty\}$ (6)
and this space is equipped with the $\|\,\cdot\,\|_{[E_{2},E_{1}]_{\theta,r}}$-norm. These spaces are again Banach spaces, see [3, Proposition 1.8 in Chapter 5]. In fact, such interpolation spaces perform a kind of interpolation of Banach spaces, as they are “in between” $E_{1}$ and $E_{2}$ in the sense that $E_{1}\hookrightarrow[E_{2},E_{1}]_{\theta,r}\hookrightarrow E_{1}$, see Lemma A.2.
@@
are bounded functionals on $E$, then $E$ is called a Banach space of functions (BSF), see [19] for more details. These spaces recently attracted interest, as they yield insights into neural networks [2].
If additionally, the norm of such a Banach space of functions $H$ is induced by an inner product, for each $x\in X$, we can find a function $k(\cdot,x)\in H$ that is associated to $\delta_{x}$ by the Riesz representation theorem. Then $H$ is a reproducing kernel Hilbert space (RKHS) on $X$ with kernel $k$. We refer to [16, Chapter 4] for fundamental results on RKHS.
@@
see e.g. [18, Lemma 2.2]. Let us denote the non-zero eigenvalues of $T_{k}$ in non-increasing order by $\mu=(\mu_{i})_{i\in I}$, where $I=\{1,...,n\}$ or $I=\mathbb{N}$. Due to the positivity and compactness of $T_{k}$, we know that $\mu_{i}>0$ and, in case that $I$ is infinite, $\lim_{i\rightarrow\infty}\mu_{i}=0$.
As proven in [18, Lemma 2.12], in such a situation, there are functions $e_{i}\in H$ such that $([e_{i}]_{\sim})_{i\in I}\subset\mathrm{L}^{2}(\nu)$ forms an ONS in $\mathrm{L}^{2}(\nu)$ consisting of eigenvectors of the integral operator $T_{k}$, such that additionally $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ forms an ONS inside $H$.
@@
Assumption H. Suppose that $(X,\mathcal{A},\mu)$ is a measure space and that $H$ is an RKHS of a measurable kernel $k$ on $X$ such that $H$ is compactly embedded into $\mathrm{L}^{2}(\nu)$. Then, for a suitable index set $I=\{1,...,n\}$ or $I=\mathbb{N}$, we denote the eigenvalues of $T_{k}:\mathrm{L}^{2}(\nu)\rightarrow\mathrm{L}^{2}(\nu)$ by $\mu_{i}$ and the associated eigenfunctions by $e_{i}$ for $i\in I$.
Assumption H+. Let Assumption H be satisfied. In addition, let the measure space $(X,\mathcal{A},\mu)$ be complete and $\mu$ be $\sigma$-finite. Furthermore, assume that $k$ is a bounded kernel.
@@
Our goal is to describe the interpolation spaces from (2) together with their norms by a spectral representation, i.e. in terms of eigenvalues and eigenfunctions of the operator $T_{k}$. To this end, for a “weight function” $w:I\rightarrow(0,\infty)$, we consider the following partition of $I$:
$M_{j}:=\left\{i\in I\,\middle|\,2^{j}<w(i)\leq 2^{j+1}\right\},\quad j\in\mathbb{Z}\text{.}$ (9)
@@
It is not hard to see that, as the $\mu_{i}$ are positive, $\|\cdot\|_{(\mu,\theta,r)}$ is in fact a norm on $\ell^{(\mu,\theta,r)}$. The choice of powers of $2$ in (9) is for notational convenience. Replacing it by any $s>1$ yields equivalent norms. It will turn out that $\ell^{(\mu,\theta,r)}$ is actually a Banach space, see Remark 5.8. Note that in the case $r=2$, one finds that $\ell^{(\mu,\theta,r)}=\ell^{2}(\mu^{-\theta})$ with equal norms. In consequence, the following definition generalises the notion of a power of an RKHS from [18, Section 4].
@@
Let $H$, $e_{i}$ and $\mu_{i}$ be as introduced in Assumption H. Then the power space of $H$ to the parameters $0<\theta<1$ and $1\leq r\leq\infty$ is defined by
$H_{\nu}^{(\theta,r)}:=\left\{f\in\mathrm{L}^{2}(\nu)\,\middle|\,\exists b\in\ell^{(\mu,\theta,r)}\text{ with }f=\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\}$ (12)
@@
$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}:=\|b\|_{(\mu,\theta,r)}\text{.}$ (13)
Note that this norm is well defined due to the uniqueness of the Fourier coefficients $b_{i}$. Recall that in [18, Section 4] only the spaces $H_{\nu}^{\theta,2}$ were introduced and investigated. Moreover, for $r\neq 2$, the spaces $H_{\nu}^{(\theta,r)}$ are no longer Hilbert spaces.
@@ 3 Results
Our three main results are provided in the following. The first result states the norm equivalence of the power spaces $H_{\nu}^{\theta,r}$ to the interpolation spaces of the form (2), as previously mentioned. This is the foundation for the subsequent result, which shows that we can view these interpolation spaces as a BSF, in case they are embedded into $\mathrm{L}^{\infty}(\nu)$. The third result then analyses in which situations such an embedding exists.
@@
Let $0<\theta<1$ and $1\leq r\leq\infty$ and suppose that $H$ is as in Assumption H. Then
$H_{\nu}^{(\theta,r)}=[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}$ (14)
@@
Let $(X,\mathcal{A},\nu)$ and $H$ be as in Assumption H+. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous. Then there is a BSF $\overline{H}_{\nu}^{(\theta,r)}$ on $X$ such that the operator
$\overline{H}_{\nu}^{(\theta,r)}\rightarrow H_{\nu}^{(\theta,r)},\quad f\mapsto[f]_{\sim}$ (15)
@@
$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (16)
The existence of an embedding of the power space $H_{\nu}^{(\theta,r)}$ into $\mathrm{L}^{\infty}(\nu)$ seems to be a crucial requirement of Theorem 3.2. Let us elaborate in which situations we can expect such an embedding. For fine-index $r=2$, this question has been approached in [18, Section 5]. We extend this approach to $1\leq r\leq\infty$.
@@
Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$, where $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Then the following statements are equivalent:
@@
    $\operatorname*{\mathrm{ess-sup}}_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa$ (17)

### v2

@@ Abstract
In statistical learning theory, interpolation spaces of the form $\mathrm{L}^{2},H]_{\theta,r}$, where $H$ is a reproducing kernel Hilbert space, are in widespread use. So far, however, they are only well understood for fine index $r=2$. We generalise existing results from $r=2$ to all possible values of $r$. In particular, we present a spectral decomposition of such spaces, analyse their embedding properties, and describe connections to the theory of Banach spaces of functions. Additionally, we present example applications of our results to regularisation error estimation in statistical learning.
Mathematics Subject Classification (2020). Primary 46E22; Secondary 46B70, 47B34, 47G10, 47B32, 47A70, 68T05.
Key Words. Reproducing Kernel Hilbert spaces, Interpolation spaces, Integral operators, Eigenvalues, Statistical learning theory.
@@ 1 Introduction
Since their introduction, reproducing kernel Hilbert spaces (RKHS) have proven to be fundamental objects in various branches of mathematical theory and in applications. They possess many desirable properties, as they combine the geometry of a Hilbert space with the continuous evaluation functionals of a function space. Apart from applications in functional analysis and numerics, statistical learning theory takes advantage of this situation by considering a large variety of kernel-based algorithms.
@@
are important for the analysis of kernel-based least squares regression, see e.g. [7], for kernel interpolation, see e.g. [16], for estimations of the conditional mean embedding, see e.g. [18], and for the analysis of learning curves, see e.g. [17]. Similarly, these spaces occur naturally for certain questions related to Gaussian processes [23] [15] [28] [24], which recently have drawn interest due to their connection to neural networks [13]. Thereby, the associated kernels are strongly related to Sobolev spaces [10], which are an important application of this paper.
The main reason for choosing $2$ as the fine index in (1) is the spectral representation of these interpolation spaces. More precisely, the spaces (1) are given as the range of powers of the integral operator associated to the RKHS [27, Section 4]. Using eigenvalues and eigenfunctions of this operator hence yields many insights into interpolation spaces and is very helpful in statistical learning theory [25] [14].
@@
  • In certain boundary cases, interpolation spaces of fine index $1$ enable embeddings into $\mathrm{L}^{\infty}(\nu)$ that are not possible for other fine indexes. For examples and applications of such embeddings, see e.g. [26, p. 4].
  • Interpolation spaces of fine index $\infty$ allow for a precise analysis of approximation errors, see e.g. [22, Theorem 3.1].
@@
that is, we allow for all possible fine indexes $1\leq r\leq\infty$. Our aim is to present methods and results that treat such interpolation spaces for all fine indexes $r$ in a joint fashion. The results we will derive address the following main questions:
@@
  • The interpolation spaces consist of equivalence classes of functions. When can we identify the spaces (2) with Banach spaces consisting of functions?
  • When are the spaces (2) continuously embedded into $\mathrm{L}^{\infty}(\nu)$?
The first question is clearly motivated by the practicality of spectral representations, as indicated by the papers mentioned above. Positive answers to the second question have applications in all situations where a point evaluation of functions is necessary. A main example is given by the theory of Gaussian processes in the context of machine learning, as paths cannot be equivalence classes but need to be defined everywhere, see e.g. [20, Assumption 1], or by the analysis of function spaces generated by neural networks, see e.g. [2]. The third question will turn out to be tightly connected to the second one, but applications go beyond that. In particular, the statistical analysis of kernel-based learning algorithms relies on such embeddings, see e.g. [25, Theorem 7.23] [7, Section 3] [26, p. 4] [19, Section 4]. Our results on the second and third question improve earlier findings such as [23, Section 5] [27, Section 5] even in the Hilbert space case $r=2$.
This work is organised as follows: we start by introducing the necessary notation and fundamental definitions in Section 2. Then the main results are provided in Section 3. In Section 4, we present an example application to error estimation in statistical learning theory. All proofs and the ideas behind the results will be given in Section 5. Finally, some auxiliary results can be found in the three appendices.
@@ 2 Preliminaries
Given normed $\mathbb{R}$-vector spaces $E_{1}$ and $E_{2}$, we use $E_{1}\simeq E_{2}$ to denote isometric isomorphism, $E_{1}\triangleq E_{2}$ for norm equivalent spaces and $E_{1}\hookrightarrow E_{2}$ if $E_{1}\subset E_{2}$ and the embedding is continuous.
For a measurable space $(X,\mathcal{A})$, we denote the set of measurable functions $X\rightarrow\mathbb{R}$ by $\mathcal{L}^{0}(X)$. The set of bounded functions in $\mathcal{L}^{0}(X)$ is denoted by $\mathcal{L}^{\infty}(X)$, and this space is equipped with the supremum norm.
@@
Let us introduce some background from the theory of interpolation spaces of the real interpolation method. For more details, see [3, Chapter 5] or [4, Chapter 3]. In the following, we only consider interpolation spaces of Banach spaces $E_{1}$ and $E_{2}$ where $E_{2}\hookrightarrow E_{1}$.
To define the real interpolation spaces, we use the so-called $K$-method. To this end, for all $x\in E_{1}$ and $t>0$, we define the $K$-functional as
$K(x,t,E_{1},E_{2}):=\inf_{y\in E_{2}}\left(\|x-y\|_{E_{1}}+t\|y\|_{E_{2}}\right)\text{.}$ (3)
For $x\in E_{1}$ and the parameters $0<\theta<1$ and $1\leq r<\infty$, this enables us to set
$\|x\|_{[E_{1},E_{2}]_{\theta,r}}:=\left(\int_{0}^{\infty}t^{-\theta r-1}K^{r}(x,t,E_{1},E_{2})\mathrm{d}t\right)^{\frac{1}{r}}$ (4)
@@
$\|x\|_{[E_{1},E_{2}]_{\theta,\infty}}:=\sup_{t>0}\left(t^{-\theta}K(x,t,E_{1},E_{2})\right)\text{.}$ (5)
@@
$[E_{1},E_{2}]_{\theta,r}:=\{x\in E_{1}\,:\,\|x\|_{[E_{1},E_{2}]_{\theta,r}}<\infty\}$ (6)
and this space is equipped with the $\|\,\cdot\,\|_{[E_{1},E_{2}]_{\theta,r}}$-norm. These spaces are again Banach spaces, see [3, Proposition 1.8 in Chapter 5]. In fact, such interpolation spaces perform a kind of interpolation of Banach spaces, as they are “in between” $E_{1}$ and $E_{2}$ in the sense that $E_{2}\hookrightarrow[E_{1},E_{2}]_{\theta,r}\hookrightarrow E_{1}$, see Lemma A.2.
@@
are bounded functionals on $E$, then $E$ is called a Banach space of functions (BSF), see [29] for more details.
If additionally, the norm of such a Banach space of functions $H$ is induced by an inner product, for each $x\in X$, we can find a function $k(\cdot,x)\in H$ that is associated to $\delta_{x}$ by the Riesz representation theorem. Then $H$ is a reproducing kernel Hilbert space (RKHS) on $X$ with kernel $k$. We refer to [25, Chapter 4] for fundamental results on RKHS.
@@
see e.g. [27, Lemma 2.2]. Let us denote the non-zero eigenvalues of $T_{k}$ in non-increasing order by $\mu=(\mu_{i})_{i\in I}$, where $I=\{1,...,n\}$ or $I=\mathbb{N}$. Due to the positivity and compactness of $T_{k}$, we know that $\mu_{i}>0$ and, in case that $I$ is infinite, $\lim_{i\rightarrow\infty}\mu_{i}=0$.
As proven in [27, Lemma 2.12], in such a situation, there are functions $e_{i}\in H$ such that $([e_{i}]_{\sim})_{i\in I}\subset\mathrm{L}^{2}(\nu)$ forms an ONS in $\mathrm{L}^{2}(\nu)$ consisting of eigenvectors of the integral operator $T_{k}$, such that additionally, $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ forms an ONS inside $H$.
@@
Assumption H. Suppose that $(X,\mathcal{A},\nu)$ is a measure space and that $H$ is an RKHS of a measurable kernel $k$ on $X$ such that $H$ is compactly embedded into $\mathrm{L}^{2}(\nu)$. Then, for a suitable index set $I=\{1,...,n\}$ or $I=\mathbb{N}$, we denote the eigenvalues of $T_{k}:\mathrm{L}^{2}(\nu)\rightarrow\mathrm{L}^{2}(\nu)$ by $\mu_{i}$ and the associated eigenfunctions by $e_{i}$ for $i\in I$.
Assumption H+. Let Assumption H be satisfied. In addition, let $\nu$ be $\sigma$-finite. Furthermore, assume that $k$ is a bounded kernel.
Let us recall two concepts from topology and measure theory. Firstly, let $X$ be a set and suppose that $M$ is a set of functions $X\rightarrow\mathbb{R}$. Then the coarsest topology $\tau(M)$ on $X$ such that all functions in $M$ are continuous is called the initial topology on $X$ with respect to $M$. Secondly, let $(X,\mathcal{A})$ be a measurable space and $(X,\mathcal{T})$ a topological space. Then a measure $\nu$ on $(X,\mathcal{A})$ is called $\mathcal{T}$-positive, if $\mathcal{T}\subset\mathcal{A}$, i.e. each open set is measurable, and if for all non-empty open sets $A$, it holds that $\nu(A)>0$.
Assumption T. Let $(X,\mathcal{A},\nu)$ be a measure space such that $\mathrm{L}^{2}(\nu)$ is separable. Moreover, let $H$ be an RKHS over $X$ of a measurable kernel $k$ such that Assumption H is satisfied and $\nu$ is $\tau(H)$-positive. We use $I$, $e_{i}$ and $\mu_{i}$ as in Assumption H.
Remark 2.1.
We give a sufficient condition that allows to verify Assumption T in the most common situations. To this end, suppose a topology $\mathcal{T}$ on $X$ is given such that
  1. (i)
    $\mathcal{T}\subset\mathcal{A}$, i.e. $\mathcal{A}$ contains the Borel-$\sigma$-algebra of $\mathcal{T}$,
  2. (ii)
    $\nu$ is $\mathcal{T}$-positive, and
  3. (iii)
    $k$ is $\mathcal{T}\otimes\mathcal{T}$-continuous.
Then, due to (iii), all functions in $H$ are $\mathcal{T}$-continuous and hence $\tau(H)\subset\mathcal{T}$. By (ii), it follows that all non-empty open sets in $\tau(H)$ have positive measure, so $\nu$ is $\tau(H)$-positive as well. Hence $\tau(H)\subset\mathcal{A}$ is $\nu$-positive. Note however that the separability of $\mathrm{L}^{2}(\nu)$ has to be checked separately.
In particular, assume that $X$ is a subset of $\mathbb{R}^{n}$ that satisfies $\mathrm{clos}(\mathrm{int}(X))=\mathrm{clos}(X)$, where $\mathrm{clos}$ and $\mathrm{int}$ denote the closure and the interior, respectively. Then Assumption T is satisfied if we use the Borel-$\sigma$-algebra and the Lebesgue measure with a continuous kernel on $X$. More generally, we may replace $\mathbb{R}^{n}$ by a second-countable locally compact group equipped with the Borel-$\sigma$-algebra and the Haar measure.
@@
Our goal is to describe the interpolation spaces from (2) together with their norms by a spectral representation, i.e. in terms of eigenvalues and eigenfunctions of the operator $T_{k}$. To this end, for a “weight sequence” $\mu:I\rightarrow(0,\infty)$, we consider the following partition of $I$:
$M_{j}:=\left\{i\in I\,\middle|\,2^{j}<\mu_{i}\leq 2^{j+1}\right\},\quad j\in\mathbb{Z}\text{.}$ (9)
@@
It is not hard to see that, as the $\mu_{i}$ are positive, $\|\cdot\|_{(\mu,\theta,r)}$ is in fact a norm on $\ell^{(\mu,\theta,r)}$. The choice of powers of $2$ in (9) is for notational convenience. Replacing it by any $s>1$ yields equivalent norms. It will turn out that $\ell^{(\mu,\theta,r)}$ is actually a Banach space, see Remark 5.8. Note that in the case $r=2$, one finds that $\ell^{(\mu,\theta,r)}=\ell^{2}(\mu^{-\theta})$ with equal norms. In consequence, the following definition generalises the notion of a power of an RKHS from [27, Section 4].
@@
Let $H$, $e_{i}$ and $\mu_{i}$ be as introduced in Assumption H. Then the power space of $H$ to the parameters $0<\theta<1$ and $1\leq r\leq\infty$ is defined as
$[H]_{\sim}^{(\theta,r)}:=\left\{f\in\mathrm{L}^{2}(\nu)\,\middle|\,\exists b\in\ell^{(\mu,\theta,r)}\text{ with }f=\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\}$ (12)
@@
$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}:=\|b\|_{(\mu,\theta,r)}\text{.}$ (13)
Note that this norm is well-defined due to the uniqueness of the Fourier coefficients $b_{i}$. Recall that in [27, Section 4], only the spaces $H]_{\sim}^{(\theta,2)}$ were introduced and investigated. Moreover, for $r\neq 2$, the spaces $H]_{\sim}^{(\theta,r)}$ are no longer Hilbert spaces.
@@ 3 Results
Our main results are provided in the following. The first result states the norm equivalence of the power spaces $H]_{\sim}^{\theta,r}$ to the interpolation spaces of the form (2), as previously mentioned. This is the foundation for the subsequent two results, which show that we can view these interpolation spaces as BSFs, in case they are embedded into $\mathrm{L}^{\infty}(\nu)$. The last result then analyses in which situations such an embedding exists.
@@
Let $0<\theta<1$, $1\leq r\leq\infty$, and suppose that $H$ is as in Assumption H. Then
$[H]_{\sim}^{(\theta,r)}=[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}$ (14)
@@
Let $(X,\mathcal{A},\nu)$ and $H$ be as in Assumption H+. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous. Then there is a BSF $\overline{H}_{\nu}^{(\theta,r)}$ on $X$ such that the operator
$\overline{H}_{\nu}^{(\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)},\quad f\mapsto[f]_{\sim}$ (15)
@@
$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (16)
Additionally, there is a $\nu$-null set $N\subset X$ such that for all $f\in\overline{H}_{\nu}^{(\theta,r)}$, there is a $(b_{i})_{i\in I}\in\ell^{(\mu,\theta,r)}$ such that $f(x)=\sum_{i\in I}b_{i}e_{i}(x)$ for all $x\in X\setminus N$.
Theorem 3.2 ensures the existence of a BSF, whose functions may differ from the actual functions in $H$ on sets of measure zero. Under Assumption T, we can make the BSF more explicit. Then we can take orthogonal series of eigenfunctions as elements of the BSF, which finally enables to embed $H$ into this BSF, as the following Theorem states.
@@
Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption T. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous. Then the following statements are true:
  1. (i)
    The series $\sum_{i\in I}b_{i}e_{i}$ converges uniformly for all $b\in\ell^{(\mu,\theta,r)}$.
  2. (ii)
    The space
    ${H}_{\nu}^{(\theta,r)}:=\left\{\sum_{i\in I}b_{i}e_{i}\,\middle|\,b\in\ell^{(\mu,\theta,r)}\right\}$ (17)
  3. (iii)
    The mapping
    ${H}_{\nu}^{(\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)},\quad\sum_{i\in I}b_{i}e_{i}\mapsto\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ (18)
    is an isometric isomorphism.
  4. (iv)
    It holds that
    $\|{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{C}_{b}(X,\tau(H))\|=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (19)
  5. (v)
    We have $H\hookrightarrow{H}_{\nu}^{(\theta,r)}$.
The existence of an embedding of the power space $H]_{\sim}^{(\theta,r)}$ into $\mathrm{L}^{\infty}(\nu)$ seems to be a crucial requirement of Theorems 3.2 and 3.3. Let us elaborate in which situations we can expect such an embedding. For fine index $r=2$, this question has been approached in [27, Section 5]. We extend this approach to $1\leq r\leq\infty$.
Theorem 3.4.
Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H. Moreover, let $\kappa>0$, $0<\theta<1$, and $1\leq r\leq\infty$, where $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Consider the following statements:
@@
    $\sup_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa\text{.}$ (20)